OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4 ? a). y = 4x - 11 b). y = 4x + 11 c). y = 4x + 7 d). y = 4x - 7

OpenStudy (anonymous):

y - y1 = m(x - x1) using slope(m) = 4 and y1 = -1 and x1 = -3

OpenStudy (anonymous):

now you sub all your stuff in....can you do that ?

OpenStudy (anonymous):

not really :/

OpenStudy (anonymous):

y - y1 = m(x - x1) y - (-1) = 4(x - (-3) y + 1 = 4(x + 3) y + 1 = 4x + 12 y = 4x + 12 -1 y = 4x + 11 do you understand ?

OpenStudy (anonymous):

yes i think i get it , so basically i just plug in everything to get my answer . thank you .

OpenStudy (anonymous):

your welcome :)